part 1- A mass of 5KG is on top of a building of heights 10M calculate its potential energy the mass then falls at a speed of 8M/S in 4 seconds find its KE and momentum P what is the force and impulse the mass transfers to the ground

part-2- For the above problem find the work done by gravity if the mass falls a distance of 2 meters if the mass falls for 11 seconds what is the energy expanded by this force?
1
Expert's answer
2019-11-22T10:57:26-0500

part 1


Potential energy:


"E=mgh=(5)(9.8)(10)=490\ J"


"KE=0.5mv^2=0.5(5)(8)^2=160\ J"

 

"P=mv=(5)(8)=40\frac{kgm}{s}"

​ The impulse the mass transfers to the ground:


"I=P=40\frac{kgm}{s}"

 The force:


"F=\frac{I}{t}=\frac{40}{4}=10\ N"

part 2



"W=mgh=(5)(9.8)(2)=98\ J"

"E=mgh'=0.5mg^2t^2"

"E=0.5(5)((11)(9.8))^2=29000\ J=29\ kJ"